A negatively charged particle, with charge q = 5.10×10−6 C, has a velocity of 416 m/s in the positive x-direction, moves into a region with a magnetic field and an electric field. The magnetic field, has a magnitude of 1.50 T, and is pointing in the positive y-direction. The electric field, has a magnitude of 4.00×103 N/C, and points in the positive z-direction. What is the value of the net force on the charged particle?

A. 2.36×10−2 N, negative z-direction

B. −1.72×10−2 N, negative z-direction

C. 1.72×10−2 N, positive z-direction

D. 12.7 N, positive z-direction
Answer #1

Force on Charged particle in magnetic field and Electric field simultaneously = Q ( v× B+ E)

= 5.1× 10-6(416× 1.5 + 4× 103)

=2.36 × 10-2 N

And direction = negative z direction

Thus it's answer is option number a.